OpenStudy (anonymous):

Solve the system by the method of substitution. 3y = –1/2x+ 2 y = –x + 9 (3, 6) (20, –4) (10, –1) (–1, 8)

OpenStudy (anonymous):

Will give medal

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@asnaseer

OpenStudy (anonymous):

Take each set of points and plug them into one of the equations.

OpenStudy (anonymous):

if i do that how do i know which one is right

HanAkoSolo (jamierox4ev3r):

because it is clearly stated that y=-x+9, you can plug it in to make one big equation

HanAkoSolo (jamierox4ev3r):

3(-x+9)=-1/2x+2

OpenStudy (amistre64):

in order to BE a solution to the system of equations .... the correct option has to make both cases true

HanAkoSolo (jamierox4ev3r):

-3x+27=-1/2x+2 +3x -2 +3x -2 25= 5/2x x=10

OpenStudy (anonymous):

kk thx

HanAkoSolo (jamierox4ev3r):

can you solve for the y-variable @isiah ?

OpenStudy (anonymous):

i believe i can if i cant can i ask you for help

HanAkoSolo (jamierox4ev3r):

aweseom just hit me up \(\Huge\ddot\smile\)